# On the Performance Analysis of Momentum Method: A Frequency Domain Perspective

### v1

@@ 3.2 Unorthodox Momentum Systems
Figure 3: (Up) Analysis of the (dynamic) magnitude responses in the early and late training stages for Standard-SGDM with low-pass gain momentum defined in Equation (11). The solid lines denote the magnitude responses in the early stages, and the dashed lines denote the magnitude responses in the late stages. (Down) The comparison between the gradient norms and momentum norms for Standard-SGDM with low-pass gain momentum. Left Column: increasing sequence. Middle Column: fixed sequence. Right Column: decreasing sequence.
@@
Example 3: Low-Pass Gain Momentum. In low-pass gain momentum, the system transits from an all-pass to a narrower low-pass gain filter as $u_{t}$ increases, amplifying low-frequency components while attenuating high-frequency components. Fig. 3(a) - 3(c) show the corresponding dynamic magnitude responses over time. A large $\mu$ corresponds to the slow shifts. Decreasing $u_{t}$ reverses the trend, heavily amplifying low-frequency components early and relaxing this effect over time. Fixed $u_{t}$ maintains constant filters, in which larger $u_{t}$ amplifies low-frequency components more aggressively. Fig. 3(d) - 3(f) demonstrate larger momentum norms compared to gradient norms, indicating the amplification effects in gain filters. Larger $u_{t},\nu$ and smaller $\mu$ lead to more reduced momentum norms, which validates the time-variant filtering characteristics of orthodox momentum systems. Test results in Table 2 indicate that increasing $u_{t}$ with appropriate $\mu$ outperforms the scenarios using fixed and decreasing sequences of $u_{t}$. We also find that smaller $\nu$ yields worse accuracy in test sets. From these results, we conclude that amplifying low-frequency gradient components and attenuating high-frequency ones in a proper way, improves test set performance.
Example 4: High-Pass Gain Momentum. High-pass gain momentum mirrors the dynamic magnitude response behavior of low-pass gain systems (see Fig. 7 in App B.2). Increasing $u_{t}$ gradually amplifies high-frequency gradient components and attenuates low-frequency ones. Decreasing $u_{t}$ reverses this pattern, heavily amplifying high-frequency components early on. Fixed constructions more aggressively amplify high-frequency components for larger $u_{t}$. The comparison of gradient norms and momentum norms can be found in Appendix B.2. Test accuracy in Table 2 shows that fixed constructions with larger $u_{t}$ and decreasing $u_{t}$ with larger $\nu$ perform worse. These findings confirm that amplifying high-frequency gradients in training might be undesirable.
@@ 3.3 Discussion
The differences in norm comparison and test accuracy between orthodox and unorthodox momentum systems validate the distinction between EMA-SGDM and Standard-SGDM from the signal processing perspective. EMA-SGDM only possesses attenuating filter effects, while Standard-SGDM can, at the same time, amplify and attenuate different frequency gradient components. Moreover, we show that Standard-SGDM with proper momentum coefficients can consistently outperform EMA-SGDM. These results clarify the difference between coupled and decoupled momentum coefficients and indicate that decoupling the momentum coefficients is usually beneficial, which answers Question 1.
Regarding Question 2, the test results show that decoupled momentum coefficients with a properly increasing $u_{t}$ and fixed $v_{t}$ can achieve better performance. In particular, our empirical findings reveal the following insights in Convolutional Neural Network (CNN) training: 1) high-frequency gradient components are undesired in the late stages of training; 2) preserving the original gradient in the early stages leads to improved test set accuracy; 3) gradually amplifying low-frequency gradient components enhance the generalization performance. Furthermore, we find that these insights are also adaptable in various learning areas (see Section 5). Based on these insights, it may be possible to design a more effective optimizer by appropriately adjusting the momentum coefficients.
@@ 4 Frequency-based Optimizer
As suggested by our frequency domain analysis framework, achieving better test performance is equivalent to finding an appropriate dynamic filter-changing pattern for momentum systems. Based on this idea, we propose FSGDM, a heuristic optimizer that dynamically adjusts momentum filtering characteristics. Furthermore, in order to explore the potential optimal strategies of our proposed FSGDM based on the findings in Section 3.3, several sets of experiments in various deep-learning tasks are conducted.
@@ 4.1 Frequency Stochastic Gradient Descent with Momentum
The proof of Proposition 4.1 is deferred in Appendix A.3. By this, we show that the dynamic magnitude response of a well-performed FSGDM can be adaptable to various tasks. In the following subsection, we explore the optimal scaling factor $c$ and momentum coefficient $v$ for FSGDM.
@@ 4.2 Empirical Exploration of Optimal Settings for FSGDM
Specifically, we empirically explore the optimal parameter selection of FSGDM across three different image classification tasks by first sweeping $c$ and $v$ within the ranges of $(0,1)$ and $[0.5,3]$, respectively. Specifically, we conduct three sets of experiments using the same codebase (See Appendix C for more training details): (1) training ResNet18 for 100 epochs on CIFAR-10, (2) training ResNet34 for 100 epochs on Tiny-ImageNet (Le and Yang,, 2015), and (3) training ResNet50 for 300 epochs on CIFAR-100. We also explore the optimal parameter selection on one Natural Language Processing (NLP) task in Appendix B.7. By finding the parameter selections with better test performance in different tasks, we try to empirically summarize the law of optimal parameter selection.
@@
The results in Fig. 4 show that there exists an optimal zone where relatively better test accuracy results can be achieved. When the momentum coefficient $v$ is fixed, the test accuracy shows an initial increase followed by a decline as the scaling factor $c$ increases. In Appendix B.8, we plot the magnitude responses and the test accuracy results of the black points in Fig. 4 and find that these parameter selections have similar dynamic magnitude responses and test accuracy curves. Thus, we suspect that the parameter selections with similar dynamic magnitude responses will lead to close performance. More discussions are in Appendix B.8.
@@ 5 Experiments
To verify the generalization of the proposed FSGDM, we perform a large-scale comparison across vision classification tasks, natural language processing tasks and reinforcement learning (RL) tasks. We compare the test performance of FSGDM and conventional SGD-based momentum optimizers, including Standard-SGDM and EMA-SGDM. We set $u_{t}=0.9,v_{t}=1$ for Standard-SGDM, and $u_{t}=0.9$ for EMA-SGDM, which are the common momentum coefficient selections in training neural networks. For a fair comparison and convenience, we set $c=0.033,v=1$, which is one of the black points in the optimal zone in Fig. 4, for FSGDM. Note that other combinations of $c$ and $v$ in the optimal zone can also be selected. For the other adjustable parameters in Algorithm 1, we set N to 300 as mentioned at the end of Section 2.1, and set $\Sigma$ as the number of total training steps. Notably, since our focus is on comparing the performance of different optimizers, we do not fine-tune every parameter for each individual model but use the same hyperparameters across all models for convenience. See Appendix C for more experimental details.
@@
CNNs on Image Classification. We perform four sets of experiments with different datasets in computer vision tasks and use various CNN architectures for training them. Specifically, we select: (a) VGG16 and ResNet18 for CIFAR-10; (b) ResNet50 and DenseNet121 (Huang et al.,, 2017) for CIFAR-100; (c) ResNet34 and MobileNet (Howard,, 2017) for Tiny-ImageNet; (d) ResNet50 for ILSVRC 2012 ImageNet Russakovsky et al., (2015). For each task, we report the mean and standard error (as subscripts) of test accuracy for 3 runs with random seeds from 0-2. The results in Table 3 show that our FSGDM consistently achieves better test set performance. Additionally, we can observe that Standard-SGDM steadily outperforms EMA-SGDM, which aligns with our discoveries in Section 3.3.
Natural Language Processing (NLP). We conduct experiments on the IWSLT14 German-English translation task (Cettolo et al.,, 2014) to represent NLP tasks, a widely used benchmark in the community. Specifically, we train six different models encompassing a variety of architectures: two convolution-based models, FConv (Gehring et al.,, 2017) and LightConv (Wu et al.,, 2019); two LSTM-based models, vanilla LSTM (Hochreiter et al.,, 1997) and an LSTM variant, LSTM-W (Wiseman and Rush,, 2016); and two Transformer-based models (Vaswani et al.,, 2017) with different sizes, Transformer-tiny and Transformer. Model performance is reported using BLEU scores, where higher scores indicate better performance, and we summarize all results in Table 4. Compared with the baseline optimizers, the proposed FSGDM outperforms all others in this task across six different models. This shows the effectiveness of our optimizer in improving translation quality. Moreover, the consistent improvement highlights the robustness of FSGDM and its ability to generalize across different neural network structures in natural language processing tasks.
@@
Figure 5: The reward curves of EMA-SGD, Standard-SGD and FSGDM on three MuJoCo tasks.
Reinforcement Learning (RL). We evaluate FSGDM on PPO (Schulman et al.,, 2017), one of the most popular policy gradient methods in reinforcement learning. We replace the default Adam optimizer (Kingma and Ba,, 2014) in PPO with FSGDM, Standard-SGDM and EMA-SGDM. We test the three optimizers on Walked2d-v4, HalfCheetah-v4 and Ant-V4, which are actually continuous control environments simulated by the standard and widely-used engine, MuJoCo (Todorov et al.,, 2012). Following standard evaluation, we run each game under 10 random seeds (range from 0-9), and test the performance for 10 episodes every 30,000 steps. All experiments are conducted using the Tianshou codebase (Weng et al.,, 2022), a widely known RL framework. Fig. 5 presents the results on three tasks, where the solid line represents the average episode rewards during evaluation, and the shaded region indicates the 75% confidence interval. It can be easily observed that on three test games, our FSGDM achieves the highest rewards than Standard-SGDM and EMA-SGDM.
6 Conclusion
This paper proposes a frequency domain analysis framework for the momentum method. Based on the proposed framework, we find that different selections of momentum coefficients correspond to different filter characteristics of the momentum methods. The generalization performance will be significantly distinct under different time-variant momentum coefficients. Furthermore, we develop a heuristic optimizer named FSGDM which outperforms the conventional SGD-based momentum optimizers in various learning tasks. Future work may investigate how we can achieve the best filtering strategy for all the general scenarios in a more theoretical way. Moreover, extending the frequency domain analysis framework to other second-moment optimizers like RMSprop (Tieleman and Hinton,, 2012) and Adam is also an interesting topic.
@@ Acknowledgments
We would like to especially thank K. C. Ho for many enlightening discussions. The work of Xianliang Li and Sheng Xu was supported by the National Natural Science Foundation of China (62273327) and Shenzhen Science and Technology Program (KCXFZ20211020165003005). The work of Linlong Wu was supported by Luxembourg FNR CORE METSA project C22/IS/17391632.
Appendix
Appendix A Additional Derivation and Proof
A.1 Derivation of Equation (8)
@@ A.2 Derivation of Equation (9)
Appendix B Additional Experiments
In this section, we present several supplementary experiments. The detailed experimental settings are shown in Appendix C.
B.1 Dynamic Sequence Construction
There are infinite increasing or decreasing sequences. In this part, we compare the test set performance of the sequence mentioned in Equ. 4 with four other dynamic sequences. Specifically, we compare with the following four dynamic increasing sequences within Algorithm 1:
@@ B.1 Dynamic Sequence Construction
where $a_{1}$ to $a_{4}$ are scaling coefficients. For a fair comparison, we adjust the coefficients to keep the $u_{t}$ of all sequences unchanged in the beginning and ending stages. To make other types of sequences unique, we keep the $u_{t}$ of different dynamic sequences nearly unchanged in the beginning and ending stages. Table 5 displays their test accuracy results after $300$ epochs of training on CIFAR-100 using ResNet50. We ran each experiment under 3 different random seeds (0, 1, 2). Clearly, the dynamic sequence we use in Equ. 4 shows its superiority over other constructions.
@@
B.2 Additional Figures of High-pass Momentum Systems on CIFAR-100
This subsection provides the figures of the dynamic magnitude responses and norm of high-pass (gain) momentum systems mentioned in Section 3. Fig. 6 and Fig. 7 show the magnitude responses and norm comparisons of high-pass and high-pass gain momentum systems, respectively. The high-pass (gain) momentum systems preserve or even amplify rapidly fluctuating gradient components, leading to sharp oscillations in gradient norm curves and momentum norm curves across iterations.
@@ B.2 Additional Figures of High-pass Momentum Systems on CIFAR-100
Figure 6: (Up) Analysis of the (dynamic) magnitude responses in the early and late training stages for EMA-SGDM with high-pass momentum defined in Equation (10). The solid lines denote the magnitude responses in the early stages, and the dashed lines denote the magnitude responses in the late stages. (Down) The comparison between the gradient norms and momentum norms for EMA-SGDM with high-pass momentum. Left Column: increasing sequence. Middle Column: fixed sequence. Right Column: decreasing sequence.
@@
Figure 7: (Up) Analysis of the (dynamic) magnitude responses in the early and late training stages for Standard-SGDM with high-pass gain momentum defined in Equation (11). The solid lines denote the magnitude responses in the early stages, and the dashed lines denote the magnitude responses in the late stages. (Down) The comparison between the gradient norms and momentum norms for Standard-SGDM with high-pass gain momentum. Left Column: increasing sequence. Middle Column: fixed sequence. Right Column: decreasing sequence.
B.3 Additional Experiments of VGG16 on CIFAR-10
In this subsection, we provide experiments of training VGG16 on CIFAR-10. The experimental settings follow Section 3 and Appendix C. From the test accuracy in Table 6 and Table 7, we observe that the test performances and norm comparisons in different momentum methods in training VGG16 on CIFAR-10 are similar to those in training ResNet50 on CIFAR-100. This similarity implies that the empirical findings in Section 3 are applicable to various CNNs.

### v3

@@ 3.2 Unorthodox Momentum Systems
Figure 3: (Up) Analysis of the (dynamic) magnitude responses in the early and late training stages for Standard-SGDM with low-pass gain momentum defined in Equation 10. The solid lines denote the magnitude responses in the early stages, and the dashed lines denote the magnitude responses in the late stages. (Down) The comparison between the gradient norms and momentum norms for Standard-SGDM with low-pass gain momentum. Left Column: increasing sequence. Middle Column: fixed sequence. Right Column: decreasing sequence.
@@
Example 3: Low-Pass Gain Momentum. In low-pass gain momentum, the system transits from an all-pass to a narrower low-pass gain filter as $u_{t}$ increases, amplifying low-frequency components while attenuating high-frequency components. Figure 3(a) - 3(c) show the corresponding dynamic magnitude responses over time. A large $\mu$ corresponds to the slow shifts. Decreasing $u_{t}$ reverses the trend, heavily amplifying low-frequency components early and relaxing this effect over time. Fixed $u_{t}$ maintains constant filters, in which larger $u_{t}$ amplifies low-frequency components more aggressively. Figure 3(d) - 3(f) demonstrate larger momentum norms compared to gradient norms, indicating the amplification effects in gain filters. Larger $u_{t},\nu$ and smaller $\mu$ lead to more reduced momentum norms, which validates the time-variant filtering characteristics of orthodox momentum systems. Test results in Table 2 indicate that increasing $u_{t}$ with appropriate $\mu$ outperforms the scenarios using fixed and decreasing sequences of $u_{t}$. We also find that smaller $\nu$ yields worse accuracy in test sets. From these results, we conclude that amplifying low-frequency gradient components and properly attenuating high-frequency ones, improves test set performance.
Example 4: High-Pass Gain Momentum. High-pass gain momentum mirrors the dynamic magnitude response behavior of low-pass gain systems (see Figure 7 in App C.2). Increasing $u_{t}$ gradually amplifies high-frequency gradient components and attenuates low-frequency ones. Decreasing $u_{t}$ reverses this pattern, heavily amplifying high-frequency components early on. Fixed constructions more aggressively amplify high-frequency components for larger $u_{t}$. The comparison of gradient norms and momentum norms can be found in Appendix C.2. Test accuracy in Table 2 shows that fixed constructions with larger $u_{t}$ and decreasing $u_{t}$ with larger $\nu$ perform worse. These findings confirm that amplifying high-frequency gradients in training might be undesirable.
@@ 3.3 Discussion
The differences in norm comparison and test accuracy between orthodox and unorthodox momentum systems validate the distinctions between EMA-SGDM and Standard-SGDM. While EMA-SGDM possesses attenuating filter effects, Standard-SGDM can both amplify and attenuate different frequency gradient components. Moreover, our findings indicate that with appropriate momentum coefficients, Standard-SGDM consistently outperforms EMA-SGDM, showing the advantages of decoupling momentum coefficients, which answers Question 1.
Regarding Question 2, the test results show that decoupled momentum coefficients with a properly increasing $u_{t}$ and fixed $v_{t}$ can achieve better performance. In particular, our empirical findings reveal the following insights in training convolutional neural networks (CNNs): (1) high-frequency gradient components are undesired in the late stages of training; (2) preserving the original gradient in the early stages leads to improved test set accuracy; (3) gradually amplifying low-frequency gradient components enhances performance. Furthermore, we find that these insights are also adaptable in various learning areas (see Section 5). Based on these insights, it may be possible to design a more effective optimizer by appropriately adjusting the momentum coefficients.
@@ 4 Frequency-based Optimizer
As suggested by our frequency domain analysis framework, achieving better test performance is equivalent to finding an appropriate dynamic filter-changing pattern for momentum systems. Based on this idea, we propose FSGDM, a heuristic optimizer that dynamically adjusts momentum filtering characteristics. Furthermore, to explore the potential optimal strategies of our proposed FSGDM based on the findings in Section 3.3, several sets of experiments in various deep-learning tasks are conducted.
@@ 4.1 Frequency Stochastic Gradient Descent with Momentum
The proof of Proposition 1 is deferred in Appendix B.3. By this, we show that the dynamic magnitude response of a well-performed FSGDM can be adaptable to various tasks. In the following subsection, we explore the optimal scaling factor $c$ and momentum coefficient $v$ for FSGDM.
@@ 4.2 Empirical Exploration of Optimal Settings for FSGDM
Specifically, we empirically explore the optimal parameter selection of FSGDM across three different image classification tasks by first sweeping $c$ and $v$ within the ranges of $(0,1)$ and $[0.5,3]$, respectively. Specifically, we conduct three sets of experiments using the same codebase (See Appendix D for more training details): (1) training ResNet18 for 100 epochs on CIFAR-10, (2) training ResNet34 for 100 epochs on Tiny-ImageNet (Le & Yang, 2015), and (3) training ResNet50 for 300 epochs on CIFAR-100. We also explore the optimal parameter selection on one natural language processing task in Appendix C.7. By finding the parameter selections with better test performance in different tasks, we try to empirically summarize the law of optimal parameter selection.
@@
The results in Figure 4 show that there exists an optimal zone where relatively better test accuracy results can be achieved. When the momentum coefficient $v$ is fixed, the test accuracy shows an initial increase followed by a decline as the scaling factor $c$ increases. In Appendix C.8, we plot the magnitude responses and the test accuracy results of the black points in Figure 4 and find that these parameter selections have similar dynamic magnitude responses and test accuracy curves. Thus, we assume the parameter selections with similar dynamic magnitude responses will lead to close performance. More discussions are in Appendix C.8.
@@ 5 Experiments
To verify the generalization of the proposed FSGDM, we perform a large-scale comparison across vision classification tasks, natural language processing (NLP) tasks, and reinforcement learning (RL) tasks. We compare the test performance of FSGDM and conventional SGD-based momentum optimizers, including Standard-SGDM and EMA-SGDM. We set $u_{t}=0.9,v_{t}=1$ for Standard-SGDM, and $u_{t}=0.9$ for EMA-SGDM, which are the common momentum coefficient selections in training neural networks. For a fair comparison and convenience, we set $c=0.033,v=1$, which is one of the black points in the optimal zone in Figure 4, for FSGDM. Note that other combinations of $c$ and $v$ in the optimal zone can also be selected. For the other adjustable parameters in Algorithm 1, we set N to 300 as mentioned at the end of Section 2.1, and set $\Sigma$ as the number of total training steps. Notably, since our focus is on comparing the performance of different optimizers, we do not fine-tune every parameter for each model but use the same hyperparameters across all models for convenience. See Appendix D for more experimental details.
@@
Image Classification. We perform four sets of experiments with different datasets in computer vision tasks and use various CNN architectures for training them. Specifically, we select: (a) VGG16 and ResNet18 for CIFAR-10; (b) ResNet50 and DenseNet121 (Huang et al., 2017) for CIFAR-100; (c) ResNet34 and MobileNet (Howard, 2017) for Tiny-ImageNet; (d) ResNet50 for ILSVRC 2012 ImageNet Russakovsky et al. (2015). For each task, we report the mean and standard error (as subscripts) of test accuracy for 3 runs with random seeds from 0-2. The results in Table 3 show that our FSGDM consistently achieves better test set performance. Additionally, we can observe that Standard-SGDM steadily outperforms EMA-SGDM, which aligns with our discoveries in Section 3.3.
Natural Language Processing. We conduct experiments on the IWSLT14 German-English translation task (Cettolo et al., 2014) to represent NLP tasks, a widely used benchmark in the community. Specifically, we train six different models encompassing a variety of architectures: two convolution-based models, FConv (Gehring et al., 2017) and LightConv (Wu et al., 2019); two LSTM-based models, vanilla LSTM (Hochreiter & Schmidhuber, 1997) and LSTM-W (Wiseman & Rush, 2016); and two Transformer-based models (Vaswani et al., 2017) with different sizes, Transformer-tiny and Transformer. Model performance is reported using BLEU scores, where higher scores indicate better performance, and we summarize all results in Table 4. Compared with the baseline optimizers, FSGDM outperforms all others in this task across six different models. This shows the effectiveness of our optimizer in improving translation quality. Moreover, the consistent improvement highlights the robustness of FSGDM and its ability to generalize across different neural network structures in natural language processing tasks.
@@
Figure 5: The reward curves of EMA-, Standard-SGDM, and FSGDM on three MuJoCo tasks.
Reinforcement Learning. We evaluate FSGDM on PPO (Schulman et al., 2017), one of the most popular policy gradient methods in reinforcement learning. We replace the default Adam optimizer (Kingma & Ba, 2014) in PPO with FSGDM, Standard-SGDM, and EMA-SGDM. We test the three optimizers on Walked2d-v4, HalfCheetah-v4, and Ant-V4, which are continuous control environments simulated by the standard and widely-used engine, MuJoCo (Todorov et al., 2012). Following standard evaluation, we run each game under 10 random seeds (range from 0-9) and test the performance for 10 episodes every 30,000 steps. All experiments are conducted using the Tianshou codebase (Weng et al., 2022), a widely known RL framework. Figure 5 presents the results on three tasks, where the solid line represents the average episode rewards during evaluation, and the shaded region indicates the 75% confidence interval. It can be easily observed that on three test games, our FSGDM achieves higher rewards than Standard-SGDM and EMA-SGDM.
6 Conclusions
This paper proposes a frequency domain analysis framework for the momentum method. Based on the proposed framework, we find that different selections of momentum coefficients correspond to different filter characteristics of the momentum methods. Performance will be significantly different under different time-variant momentum coefficients. Furthermore, we develop a heuristic optimizer named FSGDM which outperforms the conventional SGD-based momentum optimizers in various learning tasks. Future work may explore the best filtering strategy for all general scenarios and extend the frequency domain analysis framework to other optimizers such as Adam.
@@ Acknowledgments
We would like to especially thank K. C. Ho for many enlightening discussions. The work of Xianliang Li and Sheng Xu was supported by the National Natural Science Foundation of China (62273327) and the Shenzhen Science and Technology Program (KCXFZ20211020165003005). The work of Linlong Wu was supported by Luxembourg FNR CORE METSA project C22/IS/17391632.
Appendix A Phase Response
The phase response of the momentum system in the $k$-th stage can be written as,
Appendix B Additional Derivations and Proof
B.1 Derivation of Equation 8
@@ B.3 Proof of Proposition 1
Appendix C Additional Experiments
In this section, we present several supplementary experiments. The detailed experimental settings are shown in Appendix D.
C.1 Dynamic Sequence Construction
There are infinite increasing or decreasing sequences. In this part, we compare the test set performance of the sequence mentioned in Equation 4 with four other dynamic sequences. Specifically, we compare with the following four dynamic increasing sequences within Algorithm 1:
@@ C.1 Dynamic Sequence Construction
where $a_{1}$ to $a_{4}$ are scaling coefficients. For a fair comparison, we adjust the coefficients to keep the $u_{t}$ of all sequences unchanged in the beginning and ending stages. To make other types of sequences unique, we keep the $u_{t}$ of different dynamic sequences nearly unchanged in the beginning and ending stages. Table 5 displays their test accuracy results after $300$ epochs of training on CIFAR-100 using ResNet50. We ran each experiment under 3 different random seeds (0, 1, 2). Clearly, the dynamic sequence we use in Equation 4 shows its superiority over other constructions.
@@
C.2 Additional Figures of High-pass Momentum Systems on CIFAR-100
This subsection provides the figures of the dynamic magnitude responses and norm of high-pass (gain) momentum systems mentioned in Section 3. Figure 6 and Figure 7 show the magnitude responses and norm comparisons of high-pass and high-pass gain momentum systems, respectively. The high-pass (gain) momentum systems preserve or even amplify rapidly fluctuating gradient components, leading to sharp oscillations in gradient norm curves and momentum norm curves across iterations.
@@ C.2 Additional Figures of High-pass Momentum Systems on CIFAR-100
Figure 6: (Up) Analysis of the (dynamic) magnitude responses in the early and late training stages for EMA-SGDM with high-pass momentum defined in Equation 9. The solid lines denote the magnitude responses in the early stages, and the dashed lines denote the magnitude responses in the late stages. (Down) The comparison between the gradient norms and momentum norms for EMA-SGDM with high-pass momentum. Left Column: increasing sequence. Middle Column: fixed sequence. Right Column: decreasing sequence.
@@
Figure 7: (Up) Analysis of the (dynamic) magnitude responses in the early and late training stages for Standard-SGDM with high-pass gain momentum defined in Equation 10. The solid lines denote the magnitude responses in the early stages, and the dashed lines denote the magnitude responses in the late stages. (Down) The comparison between the gradient norms and momentum norms for Standard-SGDM with high-pass gain momentum. Left Column: increasing sequence. Middle Column: fixed sequence. Right Column: decreasing sequence.
C.3 Additional Experiments of VGG16 on CIFAR-10
In this subsection, we provide experiments of training VGG16 on CIFAR-10. The experimental settings follow Section 3 and Appendix D. From the test accuracy in Table 6 and Table 7, we observe that the test performances and norm comparisons in different momentum methods in training VGG16 on CIFAR-10 are similar to those in training ResNet50 on CIFAR-100. This similarity implies that the empirical findings in Section 3 are applicable to various CNNs.